The demand function for rose is Q= a - 5p and the supply function is Q= e + 2p + 0.1t, where a and e are positive constants and t is the average temperature in a month. Show how the equilibrium quantity and price vary with temperature?


1
Expert's answer
2022-03-23T18:34:03-0400

"Qd=Qs"

"a-5p=e+2p+0.1t"

"5p+2p=a-e-0.1t"

"7p=a-e-0.1t"

"P= \frac{a-e-0.1}{7}"


Equilibrium price of roses increases and quantity of roses decreases when temperature falls.